OpenStudy (anonymous):

a load taking 200A is supplied by copper and aluminium cables connected in parallel.the total length of conductor in each cable is 200m,and each conductor has a cross sectional area of 40 mm^2..the resistivity of copper and aluminium are 0.018 megaohmmeter and 0.028 megaohmmeter. calculate: a)the voltage drop of combined cables b)the current carried by each cable c)the power wasted in each cable

OpenStudy (anonymous):

COPPER CABLE RESISTANCE= (P*L)/A=0.018*10^6*200/40*10^-4=9*10^8 ohms similarly ALUMINUM RESISTANCE is= 14*10^8 ohms we calculated the resistance values and are connected in parallel hence by CURRENT DIVISION PRINCIPLE (b)the current through each cable is given as Icopper= (200*14*10^8)/(9+14)*10^8=121.74A Ialuminum=78.3 Q(a) voltage drop each cable is Vcopper= Icopper*Rcopper=1095.3*10^8 v Valuminum=1096.2*10^8v (c) power loss is given as Pcopper=Icopper^2*Rcopper Paluminum=Ialuminum^2*Raluminum these are the answers of given problem there might be mistakes in calculation but the procedure is correct please check it out. and observe that the properties of materials place crucial role in current sharing.